OpenStudy (anonymous):

Simplify: (9x^38)/(6(x^6)^3x)

OpenStudy (anonymous):

@am!rah

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

well you can't do much with this

OpenStudy (anonymous):

here's how I can simplify it

OpenStudy (anonymous):

\[\large =\frac{ 9x^{38} }{ 6(x^{6})^{3x}}\] \[\large =\frac{ 3(3)x^{38} }{ (3)(2)(x^{18x})}\] cancels 3 top and bottom \[\large =\frac{ 3x^{38} }{ (2)(x^{18x})}\]

OpenStudy (anonymous):

and I think we can simplify the exponent more

OpenStudy (anonymous):

\[\large =\frac{ 3x^{38-18x} }{ (2)}\]

OpenStudy (anonymous):

that's as simplified as I can get. Though I'm not sure if x^(-18x) has an abs value.

OpenStudy (anonymous):

@zepdrix can you verify if this is at its simplest form?

zepdrix (zepdrix):

Ehh, looks right to me :3

OpenStudy (anonymous):

Thank you hehe :D <3

zepdrix (zepdrix):

A medal?? ;( Ah I see you're up to your old tricks again!!

OpenStudy (anonymous):

So more then?